OpenStudy (anonymous):

The average of 5 consecutive even numbers A,B,C,D and E is 52. What is the product of B and E?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (debbieg):

I think what this problem wants you to recognize is that, since you have 5, consecutive even integers that average to 52, you know exactly WHAT each of those integers is. That is, notice that you can write: A=C-4 B=C-2 D=C+2 E=C+4 Then take that average: \[\Large \dfrac{ (C-4)+(C-2)+C+(C+2)+C+4) }{ 5 }=52\] If you solve that for C (or maybe you can see what I'm saying intuitively), you'll see a relationship between C and the average, 52. Once you know C, you know B and E. So just take their product. :)